Solve the following system of equations by any method you wish.
3x - 5 = y
2y = 6x - 9

Answer by mananth(16949)

3 x -1 y = 5 .............1
-6 x + 2 y = -9 .............2
Eliminate y
multiply (1)by 2
Multiply (2) by 1
6 x -2 y = 10
-6 x 2 y = -9
Add the two equations
0 x = 1
This is absurd
there is no solution to this system
The lines are parallel
write the equations in slope intercept formthe slopes are same and the y intercepts are different.